OpenStudy (anonymous):

solve using substitution method y=x+2, y-x=8

OpenStudy (anonymous):

Do you know how to substitute?

OpenStudy (anonymous):

well this is what I have done dont know if I am doing it right. (x+2) -x=8

OpenStudy (anonymous):

2x-x =8 x=8

OpenStudy (anonymous):

OK, so y = x + 2. That means that in the second equation, you can replace y with the expression x + 2. So y - x = 8 becomes (x + 2) - x = 8 which becomes 2 = 8 which is false. Hmm. Are you sure you copied the problem correctly? If so, then you might just have a false set of equations!

OpenStudy (anonymous):

Same thing here. The problem you're doing in the comments is different.

OpenStudy (anonymous):

I will look and come right back.but I believe it is y=x+2, y-x=8

OpenStudy (anonymous):

if a system has an infinite number of solutions, use set-builder notation to write the solution set. If a system has no solution, state this. Solve using the substitution method y = x + 2, y − x = 8 This is what I am reading

OpenStudy (anonymous):

So you substituted it into x+2-x=8 Which simplifies into 2=8 I was taught that this was neither a case of infinite solutions nor a case of no solutions. I would say that this is a Contradiction.

OpenStudy (anonymous):

Not sure what @ofarrows would say

OpenStudy (anonymous):

ok ty have time for another one